OpenStudy (anonymous):

find in exact ofrm the value of: cos 315

OpenStudy (goformit100):

ofrm means ?

OpenStudy (anonymous):

form

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }=\cos(315^{o})\]

OpenStudy (anonymous):

how do you remember/calculate that?

OpenStudy (anonymous):

Unit circle

OpenStudy (anonymous):

Learn it...Love it!